You want to obtain a sample to estimate a population proportion. Based on previous evidence, you believe the population proportion is approximately p = 0.15. You would like to be 98% confident that your esimate is within 2.5% of the true population proportion. How large of a sample size is required?

Hint: Video [+]

n =
Answer #1

We have :

p(1 -р) E Za/2 п

We are given :

E 2.5%= 0.025

p=0.15

=1-0.98= 0.02

Za/2Z0.02/2 = Z0.01 =2.3263

Now,

0.15(1 0.15) 0.025 2.3263

\sqrt{n} = 2.3263*\sqrt{0.15(1-0.15)}*\frac{1}{0.025}

\sqrt{n} =33.226

33.22621103.98 1104

Hence,

Required sample size = 1104